OpenStudy (anonymous):

Jim has a bag that contained 5 blue 5 red and 5 green marbles. he drew out 1 marble at a time and then placed it back into the bag.

OpenStudy (anonymous):

2/3 or 0.667

OpenStudy (anonymous):

so that is the answer 

OpenStudy (anonymous):

*thats for the first draw since the probability of getting a blue marble is 5/15 or 1/3, red is 5/15 or 1/3 and green is also 5/15 or 1/3 The probability of not getting a red means the probability of getting blue or green marble. then the probability of not getting red is 1/3+1/3=2/3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

considering your additional data then P(blue)=8/15 P(red) =4/14 P(green) = 3/15 P(blue or green) = 8/15=3/15 = 11/15 = 0.733